# SAR-KnowLIP: Towards Multimodal Foundation Models for Remote Sensing

### v1

@@ Abstract
Cross-modal artificial intelligence has garnered widespread attention in recent years, achieving significant progress in the study of natural images. However, existing methods are mostly designed for RGB imagery, leaving a significant gap in modeling synthetic aperture radar (SAR) imagery. SAR, with its all-day, all-weather imaging capabilities, plays an irreplaceable role in remote sensing scene understanding. To address this gap, this paper proposes SAR-KnowLIP, the first universal SAR multimodal foundational model, along with reusable data and evaluation baselines. Specifically: (1) This work introduces the critical yet long-overlooked attribute of geographic information into remote sensing research, constructing SAR-GEOVL-1M (the first large-scale SAR dataset with complete geographic projection properties), covering multiple satellite platforms, 120,000 images, and 135 cities. (2) Aligned structured text is generated through a hierarchical cognitive chain-of-thought (HCoT), providing more than one million multi-dimensional semantic annotations of landforms, regional functions, target attributes, and spatial relationships. (3) We design a Self-Consistent Iterative Optimization mechanism that continuously enhances cross-modal alignment through a self-supervised closed loop of contrastive, matching, and reconstruction learning on a transferable multimodal encoder. (4) A unified evaluation benchmark is established across 11 representative downstream vision and vision-language tasks, with comparisons against 14 leading foundation models, where SAR-KnowLIP demonstrates leading performance, particularly in object counting and land-cover classification. We expect that SAR-KnowLIP’s large-scale multimodal data, transferable model architecture, and comprehensive experimental benchmark will significantly advance the development of SAR multimodal baseline models. The dataset and model are publicly available at: https://github.com/yangyifremad/SARKnowLIP.
@@ 1 Introduction
Despite the remarkable progress of cross-modal artificial intelligence in natural imagery and optical remote sensing[1],[2], the applicability of existing visual–language models (VLMs) to synthetic aperture radar (SAR) imagery remains severely limited. Most prevailing approaches are built upon modeling assumptions designed for RGB imagery[3],[4],[5],[6], which fail to effectively capture the unique imaging mechanisms and semantic characteristics of SAR data. As a result, direct transfer often leads to inadequate representation and performance degradation. Consequently, dedicated multimodal models for SAR are still largely absent, constituting a critical bottleneck that constrains the advancement of cross-modal remote sensing research toward higher-level semantic understanding.
This situation contrasts sharply with the strategic role of SAR in the remote sensing ecosystem. As an active imaging modality, SAR provides all-weather, day-and-night observation capabilities and exhibits heightened sensitivity to structural and geometric properties of ground objects[7], enabling reliable information acquisition under nighttime, extreme weather, and high cloud or precipitation conditions. Owing to these characteristics, SAR has been widely applied to disaster monitoring, oceanic and polar observation, urban management, and national security[8]. Compared with optical and multispectral imagery, SAR is better suited for long-term monitoring in complex environments, and its irreplaceable value highlights the urgent need for dedicated multimodal foundation models tailored to SAR.
@@
1) Geospatial Priors Remain Underutilized: Geographic information, a core element in Earth sciences, is crucial for RS. However, most current SAR interpretation studies remain at the visual level, neglecting geographic attributes. This stems from the extension of natural image paradigms—focused mainly on texture and structure—as well as the closed nature of SAR data sources, with many studies still relying on outdated datasets like MSTAR from the 1990s[9]. As shown in Table I, our review of 16 mainstream SAR datasets across tasks such as detection, segmentation, and recognition reveals a widespread omission of geographic metadata. This absence hinders regional analysis, terrain understanding, and spatial reasoning, ultimately limiting progress toward cognitive-level SAR interpretation.
@@
To address these challenges, we present SAR-GEOVL-1M, the first large-scale SAR image-text dataset enriched with open text description and geographic metadata, along with SAR-KnowLIP, the first universal multimodal foundational model for SAR images. This work aims to move beyond the limitations of single-modality visual perception, establishing a multimodal SAR foundation model with cognitive reasoning, semantic representation, and transferable capabilities.
In terms of image data construction, SAR-GEOVL-1M contains 120,000 images from three SAR satellite platforms, spanning multiple resolutions, 135 cities, and five landform types. To mitigate semantic inconsistency caused by cross-platform resolution variation, we introduce a Spatial Resolution Consistency (SRC) slicing strategy, which aligns semantic granularity at the geographic level and ensures uniform cognitive scale for model training.
@@
In terms of multimodal alignment modeling, this paper proposes SAR-KnowLIP, a dual-tower multimodal framework employing Vision Transformer (ViT)[28] and BERT[29] as visual and language encoders. The model jointly optimizes image-text contrastive loss (ITC), matching loss (ITM), and masked language modeling loss (MLM) to construct a unified cross-modal embedding space bridging low-level perception and high-level semantics. To mitigate semantic deviations in HCoT-guided generated text, we introduce a self-consistent iterative optimization (SCIO) module that enhances alignment accuracy and stability through a closed-loop self-supervised strategy of screening, filtering, and reconstruction.
@@
  • SAR-GEOVL-1M Dataset: This dataset represents the first large-scale SAR image and text dataset with complete geographic information. It encompasses data from three types of SAR satellite platforms, 135 cities, and multi-scale typical scenes, including over 120,000 images and more than one million text descriptions. By addressing the absence of geographic attributes in SAR image interpretation research, this dataset fills a critical gap and provides a foundational data resource for SAR multimodal modeling research.
@@
  • SAR-KnowLIP: The first visual language foundational model for SAR is developed, combining contrast, matching, and reconstruction multi-task learning to establish a cross-modal representation space for vision-language collaboration. The SCIO optimization module is introduced to dynamically enhance text accuracy and improve cross-modal alignment quality through the ”screen–filter–reconstruct” self-supervised closed-loop optimization mechanism.
  • Leading Multi-Task Generalization Capability: In various typical downstream tasks, such as target classification, detection, land feature segmentation, image captioning, image-text retrieval, and visual question answering, SAR-KnowLIP demonstrates superior semantic understanding and cross-task generalization performance compared to existing RS multimodal models.
The remainder of this paper is organized as follows: Section 2 introduces the research progress in related fields; Section 3 details the construction process of the SAR-GEOVL-1M dataset and the design of the multimodal self-supervised model; Section 4 presents experimental verification and analysis across multiple RS downstream tasks; Section 5 provides a summary of the paper and outlines potential future research directions.
@@ 3 Method
3.1 SAR-GEOVL-1M Image Data
Figure 4: SAR-GEOVL-1M image data. FUSAR-SHIP: A ship detection dataset we released in 2020. FUSAR-MAP: A land feature classification dataset we released in 2021. FUSAR-AIR: An aircraft detection dataset we released in 2024. SAR-GEOVL: A large-scale SAR multimodal dataset proposed in this study.
Given that existing SAR datasets generally lack geospatial attributes, making them difficult to support complex semantic modeling and spatial reasoning, this study’s approach first addresses the data layer. Current mainstream datasets (as shown in Table I) often lack geographic metadata, resulting in a disconnect between imagery and real-world spatial locations, limiting tasks such as regional functional reasoning and multi-scale target analysis. To this end, we constructed SAR-GEOVL-1M, the first large-scale SAR dataset to fully preserve geographic information. Its core design principle is spatial scale consistency: it fuses SAR imagery from multiple platforms and resolutions, and provides each image with WGS84 projection coordinates, enabling precise spatial positioning and adaptability to multimodal tasks.
The SAR-GEOVL-1M dataset is collected from three SAR satellites with different resolutions, including Qilu-1[46], Gaofen-3[47], and Hongtu-1[48]. Qilu-1 has a resolution of 0.2 meters and operates in the Ku band, offering high-precision urban modeling capabilities. Gaofen-3 has a resolution of 1 meter and operates in the C band, which is suitable for regional target and structure extraction. Hongtu-1 has a resolution of 3 meters and operates in the X band, ideal for modeling urban area edges and terrain structures.
@@ 3.1 SAR-GEOVL-1M Image Data
Figure 5: Cities covered by the SAR-GEOVL-1M dataset.
@@
As shown in Fig. 4, the SAR-GEOVL-1M dataset presents a clear distribution of data, sample proportions, and representative examples. It is noteworthy that this dataset is built upon several task-oriented SAR datasets previously proposed by our team—namely, FUSAR-MAP[13], FUSAR-SHIP[19], and FUSAR-AIR[51]—which have undergone large-scale expansion and data restructuring. These enhancements have significantly improved the dataset’s coverage, semantic density, and adaptability to downstream tasks. The aforementioned datasets have been downloaded over 6,000 times and are widely cited and applied in the field of SAR interpretation. SAR-GEOVL-1M represents the systematic accumulation of our team’s long-term research in SAR image understanding and provides a strong data foundation for multimodal modeling in the SAR domain.
@@ 3.2 Text Generation Guided by Hierarchical Cognitive Chain-of-Thought (HCoT)
This mechanism simulates the human spatial cognitive process, moving from global to local, enabling the model to focus on local details when describing the targets, while ensuring consistency and logical coherence with the overall context. Guided by the HCoT reasoning mechanism, more than 1 million structured texts were generated, forming the language modality data of SAR-GEOVL-1M. A detailed statistical analysis and example verification are provided in Section 4.1.
3.3 SAR-KnowLIP Overall Framework
Figure 7: The overall framework of SAR-KnowLIP. SAR images and texts from SAR-GEOVL-1M are encoded by ViT and BERT, respectively, and optimized through the SCIO module. SCIO consists of three progressive stages: (1) cross-modal preliminary representation learning, (2) information screening and proofreading, and (3) context-aware reconstruction—forming a closed-loop “screen–filter–reconstruct” process that progressively enhances cross-modal alignment and semantic representation.
After introducing geographic information and scale consistency into images and constructing hierarchical cognitive semantics into text, the key challenge is achieving effective alignment and knowledge fusion of cross-modal representations at the model level. To this end, we propose SAR-KnowLIP, a universal visual language foundational model for SAR imagery. This framework, based on the core principles of ”multimodal semantic modeling, modality alignment, and self-supervised optimization,” aims to eliminate reliance on manual annotation and achieve deep semantic understanding of SAR imagery through cross-modal collaborative learning.
Specifically, SAR-KnowLIP adopts a standardized dual-encoder architecture to separately model visual and textual representations. The visual encoder employs a vision transformer, where SAR images are divided into fixed-size patches and processed through linear projection and positional encoding before being passed to the Transformer. This yields a uniform visual embedding vector: $f_{v}\in\mathbb{R}^{d}$. The text encoder is based on BERT, which tokenizes and embeds the input text, producing a corresponding embedding vector: $f_{t}\in\mathbb{R}^{d}$ through a multi-layer Transformer. These embeddings are aligned within a shared semantic space. ViT and BERT, both Transformer-based, are widely adopted in multimodal frameworks (e.g., BLIP, ALBEF, OFA[52]) due to their strong alignment capability, generalization, and downstream task compatibility, making them well-suited as foundational encoders for our model.
To achieve multimodal interaction and semantic alignment, SAR-KnowLIP introduces three collaborative self-supervision tasks to enhance cross-modal modeling capabilities from three levels: modality alignment, semantic matching, and text generation.
@@ 21 Feedback Loop: Backpropagate optimized text
Figure 8: Text data of SAR-GEOVL-1M. (a)–(b): word clouds and length statistics of object and landform task texts; (c)–(e): Influence of HCoT on text quality.
@@ 4.1 Implementation Details
SAR-KnowLIP relies on large-scale cross-modal pre-training, which requires balancing computational efficiency with model capacity while ensuring reproducibility. Accordingly, we implemented the complete training and evaluation pipeline in PyTorch and conducted all experiments on a computing node equipped with eight NVIDIA RTX 3090 GPUs. The visual encoder is initialized from a ViT model pretrained on ImageNet, while the text encoder uses BERT weights. Considering computational efficiency and model performance, the input image size is fixed at 224×224 with a batch size of 32 during pretraining. A queue of 20,000 image-text features is maintained to support large-scale contrastive learning. The alignment loss weight $\alpha$ is set to 0.4 to strengthen cross-modal semantic association.
@@
4.2 Analysis of SAR-GEOVL-1M Text
After constructing the SAR-GEOVL-1M dataset, we further conducted statistical and quality analysis of its language modalities. Unlike existing multimodal remote sensing datasets that rely on manual or templated labeling, the data text in this study was automatically generated by a large language model, necessitating a systematic assessment of its semantic richness, task relevance, and diversity. To ensure the accuracy and interpretability of the generated text, we constructed semantic descriptions based on GPT-4.1. This model performed best in SAR scene understanding, terminology usage, and semantic coherence in comparisons with mainstream language models like Gemini and Grok, and was therefore selected as the primary generation engine.

### v2

# FUSAR-KLIP: Towards Multimodal Foundation Models for Remote Sensing

@@ Abstract
Cross-modal artificial intelligence has garnered widespread attention in recent years, achieving significant progress in the study of natural images. However, existing methods are mostly designed for RGB imagery, leaving a significant gap in modeling synthetic aperture radar (SAR) imagery. SAR, with its all-day, all-weather imaging capabilities, plays an irreplaceable role in remote sensing scene understanding. To address this gap, this paper proposes FUSAR-KLIP, the first universal SAR multimodal foundational model, along with reusable data and evaluation baselines. Specifically: (1) This work introduces the critical yet long-overlooked attribute of geographic information into remote sensing research, constructing FUSAR-GEOVL-1M (the first large-scale SAR dataset with complete geographic projection properties), covering multiple satellite platforms, 120,000 images, and 135 cities. (2) Aligned structured text is generated through a hierarchical cognitive chain-of-thought (HCoT), providing more than one million multi-dimensional semantic annotations of landforms, regional functions, target attributes, and spatial relationships. (3) We design a Self-Consistent Iterative Optimization mechanism that continuously enhances cross-modal alignment through a self-supervised closed loop of contrastive, matching, and reconstruction learning on a transferable multimodal encoder. (4) A unified evaluation benchmark is established across 11 representative downstream vision and vision-language tasks, with comparisons against 14 leading foundation models, where FUSAR-KLIP demonstrates leading performance, particularly in object counting and land-cover classification. We expect that FUSAR-KLIP’s large-scale multimodal data, transferable model architecture, and comprehensive experimental benchmark will significantly advance the development of SAR multimodal baseline models. The dataset and model are publicly available at: https://github.com/yangyifremad/SARKnowLIP.
@@ 1 Introduction
Despite the remarkable progress of cross-modal artificial intelligence in natural imagery and optical remote sensing[1],[2], the applicability of existing visual–language models (VLMs) to synthetic aperture radar (SAR) imagery remains severely limited. Most prevailing approaches are built upon modeling assumptions designed for RGB imagery[3],[4],[5],[6],[7], which fail to effectively capture the unique imaging mechanisms and semantic characteristics of SAR data. As a result, direct transfer often leads to inadequate representation and performance degradation. Consequently, dedicated multimodal models for SAR are still largely absent, constituting a critical bottleneck that constrains the advancement of cross-modal remote sensing research toward higher-level semantic understanding.
This situation contrasts sharply with the strategic role of SAR in the remote sensing ecosystem. As an active imaging modality, SAR provides all-weather, day-and-night observation capabilities and exhibits heightened sensitivity to structural and geometric properties of ground objects[8], enabling reliable information acquisition under nighttime, extreme weather, and high cloud or precipitation conditions. Owing to these characteristics, SAR has been widely applied to disaster monitoring, oceanic and polar observation, urban management, and national security[9]. Compared with optical and multispectral imagery, SAR is better suited for long-term monitoring in complex environments, and its irreplaceable value highlights the urgent need for dedicated multimodal foundation models tailored to SAR[10].
@@
1) Geospatial Priors Remain Underutilized: Geographic information, a core element in Earth sciences, is crucial for RS. However, most current SAR interpretation studies remain at the visual level, neglecting geographic attributes. This stems from the extension of natural image paradigms—focused mainly on texture and structure[11]—as well as the closed nature of SAR data sources, with many studies still relying on outdated datasets like MSTAR from the 1990s[12]. As shown in Table I, our review of 16 mainstream SAR datasets across tasks such as detection, segmentation, and recognition reveals a widespread omission of geographic metadata. This absence hinders regional analysis, terrain understanding, and spatial reasoning, ultimately limiting progress toward cognitive-level SAR interpretation.
@@
To address these challenges, we present FUSAR-GEOVL-1M, the first large-scale SAR image-text dataset enriched with open text description and geographic metadata, along with FUSAR-KLIP, the first universal multimodal foundational model for SAR images. FUSAR-KLIP is a Knowledge-guided Language-Image Pre-training model that aims to move beyond the limitations of single-modality visual perception, establishing a multimodal SAR foundation model with cognitive reasoning, semantic representation, and transferable capabilities.
In terms of image data construction, FUSAR-GEOVL-1M contains 120,000 images from three SAR satellite platforms, spanning multiple resolutions, 135 cities, and five landform types. To mitigate semantic inconsistency caused by cross-platform resolution variation, we introduce a Spatial Resolution Consistency (SRC) slicing strategy, which aligns semantic granularity at the geographic level and ensures uniform cognitive scale for model training.
@@
In terms of multimodal alignment modeling, this paper proposes FUSAR-KLIP, a dual-tower multimodal framework employing Vision Transformer (ViT)[31] and BERT[32] as visual and language encoders. The model jointly optimizes image-text contrastive loss (ITC), matching loss (ITM), and masked language modeling loss (MLM) to construct a unified cross-modal embedding space bridging low-level perception and high-level semantics. To mitigate semantic deviations in HCoT-guided generated text, we introduce a self-consistent iterative optimization (SCIO) module that enhances alignment accuracy and stability through a closed-loop self-supervised strategy of screening, filtering, and reconstruction.
@@
  • FUSAR-GEOVL-1M Dataset: This dataset represents the first large-scale SAR image and text dataset with complete geographic information. It encompasses data from three types of SAR satellite platforms, 135 cities, and multi-scale typical scenes, including over 120,000 images and more than one million text descriptions. By addressing the absence of geographic attributes in SAR image interpretation research, this dataset fills a critical gap and provides a foundational data resource for SAR multimodal modeling research.
@@
  • FUSAR-KLIP(Knowledge-guided Language-Image Pre-training): The first visual language foundational model for SAR is developed, combining contrast, matching, and reconstruction multi-task learning to establish a cross-modal representation space for vision-language collaboration. The SCIO optimization module is introduced to dynamically enhance text accuracy and improve cross-modal alignment quality through the ”screen–filter–reconstruct” self-supervised closed-loop optimization mechanism.
  • Leading Multi-Task Generalization Capability: In various typical downstream tasks, such as target classification, detection, land feature segmentation, image captioning, image-text retrieval, and visual question answering, FUSAR-KLIP demonstrates superior semantic understanding and cross-task generalization performance compared to existing RS multimodal models.
The remainder of this paper is organized as follows: Section 2 introduces the research progress in related fields; Section 3 details the construction process of the FUSAR-GEOVL-1M dataset and the design of the multimodal self-supervised model; Section 4 presents experimental verification and analysis across multiple RS downstream tasks; Section 5 provides a summary of the paper and outlines potential future research directions.
@@ 3 Method
3.1 FUSAR-GEOVL-1M Image Data
Figure 4: FUSAR-GEOVL-1M image data. FUSAR-SHIP: A ship detection dataset we released in 2020. FUSAR-MAP: A land feature classification dataset we released in 2021. FUSAR-AIR: An aircraft detection dataset we released in 2024. FUSAR-GEOVL: A large-scale SAR multimodal dataset proposed in this study.
Given that existing SAR datasets generally lack geospatial attributes, making them difficult to support complex semantic modeling and spatial reasoning, this study’s approach first addresses the data layer. Current mainstream datasets (as shown in Table I) often lack geographic metadata, resulting in a disconnect between imagery and real-world spatial locations, limiting tasks such as regional functional reasoning and multi-scale target analysis. To this end, we constructed FUSAR-GEOVL-1M, the first large-scale SAR dataset to fully preserve geographic information. Its core design principle is spatial scale consistency: it fuses SAR imagery from multiple platforms and resolutions, and provides each image with WGS84 projection coordinates, enabling precise spatial positioning and adaptability to multimodal tasks.
The FUSAR-GEOVL-1M dataset is collected from three SAR satellites with different resolutions, including Qilu-1[49], Gaofen-3[50], and Hongtu-1[51]. Qilu-1 has a resolution of 0.2 meters and operates in the Ku band, offering high-precision urban modeling capabilities. Gaofen-3 has a resolution of 1 meter and operates in the C band, which is suitable for regional target and structure extraction. Hongtu-1 has a resolution of 3 meters and operates in the X band, ideal for modeling urban area edges and terrain structures.
@@ 3.1 FUSAR-GEOVL-1M Image Data
Figure 5: Cities covered by the FUSAR-GEOVL-1M dataset.
@@
As shown in Fig. 4, the FUSAR-GEOVL-1M dataset presents a clear distribution of data, sample proportions, and representative examples. It is noteworthy that this dataset is built upon several task-oriented SAR datasets previously proposed by our team—namely, FUSAR-MAP[16], FUSAR-SHIP[22], and FUSAR-AIR[54]—which have undergone large-scale expansion and data restructuring. These enhancements have significantly improved the dataset’s coverage, semantic density, and adaptability to downstream tasks. The aforementioned datasets have been downloaded over 6,000 times and are widely cited and applied in the field of SAR interpretation. FUSAR-GEOVL-1M represents the systematic accumulation of our team’s long-term research in SAR image understanding and provides a strong data foundation for multimodal modeling in the SAR domain.
@@ 3.2 Text Generation Guided by Hierarchical Cognitive Chain-of-Thought (HCoT)
This mechanism simulates the human spatial cognitive process, moving from global to local, enabling the model to focus on local details when describing the targets, while ensuring consistency and logical coherence with the overall context. Guided by the HCoT reasoning mechanism, more than 1 million structured texts were generated, forming the language modality data of FUSAR-GEOVL-1M. A detailed statistical analysis and example verification are provided in Section 4.1.
3.3 FUSAR-KLIP Overall Framework
Figure 7: The overall framework of FUSAR-KLIP. SAR images and texts from FUSAR-GEOVL-1M are encoded by ViT and BERT, respectively, and optimized through the SCIO module. SCIO consists of three progressive stages: (1) cross-modal preliminary representation learning, (2) information screening and proofreading, and (3) context-aware reconstruction—forming a closed-loop “screen–filter–reconstruct” process that progressively enhances cross-modal alignment and semantic representation.
After introducing geographic information and scale consistency into images and constructing hierarchical cognitive semantics into text, the key challenge is achieving effective alignment and knowledge fusion of cross-modal representations at the model level. To this end, we propose FUSAR-KLIP, a universal visual language foundational model for SAR imagery. This framework, based on the core principles of ”multimodal semantic modeling, modality alignment, and self-supervised optimization,” aims to eliminate reliance on manual annotation and achieve deep semantic understanding of SAR imagery through cross-modal collaborative learning.
Specifically, FUSAR-KLIP adopts a standardized dual-encoder architecture to separately model visual and textual representations. The visual encoder employs a vision transformer, where SAR images are divided into fixed-size patches and processed through linear projection and positional encoding before being passed to the Transformer. This yields a uniform visual embedding vector: $f_{v}\in\mathbb{R}^{d}$. The text encoder is based on BERT, which tokenizes and embeds the input text, producing a corresponding embedding vector: $f_{t}\in\mathbb{R}^{d}$ through a multi-layer Transformer. These embeddings are aligned within a shared semantic space. ViT and BERT, both Transformer-based, are widely adopted in multimodal frameworks (e.g., BLIP, ALBEF, OFA[55],[56]) due to their strong alignment capability, generalization, and downstream task compatibility, making them well-suited as foundational encoders for our model.
To achieve multimodal interaction and semantic alignment, FUSAR-KLIP introduces three collaborative self-supervision tasks to enhance cross-modal modeling capabilities from three levels: modality alignment, semantic matching, and text generation.
@@ 21 Feedback Loop: Backpropagate optimized text
Figure 8: Text data of FUSAR-GEOVL-1M. (a)–(b): word clouds and length statistics of object and landform task texts; (c)–(e): Influence of HCoT on text quality.
@@ 4.1 Implementation Details
FUSAR-KLIP relies on large-scale cross-modal pre-training, which requires balancing computational efficiency with model capacity while ensuring reproducibility. Accordingly, we implemented the complete training and evaluation pipeline in PyTorch and conducted all experiments on a computing node equipped with eight NVIDIA RTX 3090 GPUs. The visual encoder is initialized from a ViT model pretrained on ImageNet, while the text encoder uses BERT weights. Considering computational efficiency and model performance, the input image size is fixed at 224×224 with a batch size of 32 during pretraining. A queue of 20,000 image-text features is maintained to support large-scale contrastive learning. The alignment loss weight $\alpha$ is set to 0.4 to strengthen cross-modal semantic association.
@@
4.2 Analysis of FUSAR-GEOVL-1M Text
After constructing the FUSAR-GEOVL-1M dataset, we further conducted statistical and quality analysis of its language modalities. Unlike existing multimodal remote sensing datasets that rely on manual or templated labeling, the data text in this study was automatically generated by a large language model, necessitating a systematic assessment of its semantic richness, task relevance, and diversity. To ensure the accuracy and interpretability of the generated text, we constructed semantic descriptions based on GPT-4.1. This model performed best in SAR scene understanding, terminology usage, and semantic coherence in comparisons with mainstream language models like Gemini and Grok, and was therefore selected as the primary generation engine.